# New ultralight scalar particles and the mass-radius relation of white dwarfs - the important role of Sirius B

### v1

@@ Abstract
dark matter – astroparticle physics – white dwarfs
pubyear: 2025pagerange: New ultralight scalar particles and the mass-radius relation of white dwarfs - the important role of Sirius BA.2
@@ 1 Introduction
Most prominently, there is the problem of dark matter, for which the SM has no candidate. Historically, dark matter searches focused on heavy, GeV-scale candidates. However, after decades of null results from indirect dark matter searches and direct detection experiments, particle theorists started to broaden the mass range they considered for their models. A particularly active field looks at ultralight dark matter (ULDM), also referred to as weakly interacting sub-eV particles (WISPs), see e.g. (Preskill et al., 1983; Abbott & Sikivie, 1983; Dine & Fischler, 1983; Arvanitaki et al., 2015). Instead of hiding dark matter at high energies or masses, detection is avoided by adding very light bosonic particles characterized by minuscule interactions with the Standard Model. In the end, this can be described by adding a new species of light scalar particles to the SM.
Another theoretical challenge for the Standard Model is the strong CP problem, which describes the violation of charge conjugation and parity symmetry: While the electro-weak interaction has an order one CP phase, no violation of CP symmetry has been detected in processes involving only the strong interaction. There is no explanation for this puzzle within the Standard Model, however, a very attractive solution is a new dynamical field, the bosonic axion (Peccei & Quinn, 1977b, a; Wilczek, 1978; Weinberg, 1978), which is also a viable dark matter candidate. Particles with similar properties, albeit not solving the strong CP problem, are also common in string theory and many other higher-dimensional models (Arvanitaki et al., 2010). These particles are commonly referred to as axion-like particles (ALPs).
@@
Astrophysics offers alternative approaches to constrain the properties of these scalar particles. The best-known example is based on stellar cooling: If additional light particles couple weakly to Standard Model matter, they get produced in stars and stellar remnants. Due to the weak coupling, they can easily escape and thus provide an additional energy loss channel for the star. Thus, stellar cooling in connection with stellar evolution data can be used to constrain these couplings (Raffelt, 1996; Springmann et al., 2025). Considering, for example, QCD axions, generically, there are axion-nucleon couplings, which lead to axion production via nucleon bremsstrahlung. Requiring neutron stars not to cool faster than observed yields leading constraints on QCD axions (Buschmann et al., 2022; Springmann et al., 2024).
@@
In this work, we will focus on a different approach to constraining particle physics from stars. Recently (Hook & Huang, 2018) have shown that new scalar fields with quadratic coupling to the Standard Model can develop a classical expectation value within stars - similarly to the Higgs mechanism, but localized within the star (see Fig. 1), a phenomenon that might occur for the QCD axion (Balkin et al., 2020). In scalar tensor theories of modified gravity, a similar phenomenon is known as scalarization (Damour & Esposito-Farese, 1993; Doneva et al., 2024; Ramazanoğlu & Pretorius, 2016; Staykov et al., 2018). This scalar field profile, especially the part that leaks out of the star can have drastic effects, see (Hook & Huang, 2018; Zhang et al., 2021; Balkin et al., 2022, 2023). Moreover, the expectation value backreacts on the stellar structure and can lead to a new ground state of stellar matter or a phase transition in the stellar equation of state (EOS). This would have drastic effects on astrophysical objects and observations, see for example (Balkin et al., 2024, 2025; Gómez-Bañón et al., 2024; Kumamoto et al., 2025; Bartnick et al., 2025). Following this train of thought, in this work, we focus on two examples: a generic light scalar field coupled to electrons and leading to a first-order phase transition, and a lighter version of the QCD axion (Hook, 2018; Di Luzio et al., 2021; Banerjee et al., 2025a).
@@
We will use the observed white dwarf MRR to constrain these models of new physics. We note that there has been a substantial recent improvement with the MRR observational data. While theoretically, within the SM, this relation is well understood and routinely used in the analysis of white dwarfs, the empirical evidence has been problematic until very recently (e.g. (Koester, 1987; Vauclair et al., 1997; Provencal et al., 1998; Joyce et al., 2018; Bédard et al., 2017)). The bulk of the white dwarf population agreed with the theoretical MRR, but the shape of the relation was almost unconstrained. This has changed dramatically over the last years due to the use of very large telescopes for high-quality optical spectra and orbital elements of eclipsing binaries, well calibrated UV spectra from the Hubble telescope, and in particular, excellent distance measurements from Gaia (Gaia Collaboration et al., 2023). This has now resulted in high precision data for about 30 white dwarfs, which we use for the comparison with our theoretical models (see Table 1).
@@
In Section 2, we will outline the particle physics models and explain how they lead to a modified EOS. Section 3 details the calculation of white dwarf models, which lead to a (modified) mass-radius relation discussed in Section 4. The resulting limits on the particle physics models are shown and discussed in Section 5.
@@ 2 The equation of state with new light scalar fields
To derive constraints on new scalar fields, we must first understand how they influence the stellar structure of white dwarfs, which was first noted in (Balkin et al., 2024) for light QCD axions. As detailed in the Appendix, we will focus on the so-called negligible-gradient limit where the length scale over which the scalar field changes is much smaller than the size of the star. Consequently, the effect of the scalar fields can be fully described by a modification of the white dwarf EOS. In this section, we will give a brief overview of the particle physics model and modified EOS. More details can be found in App. A and (Balkin et al., 2024, 2025).
Figure 1: Schematic description of a scalar field profile emerging in a white dwarf. The main panel sketches the scalar field as a function of radius $r$ in and around a white dwarf. The dashed line marks the radius of the white dwarf $r=R_{\mathrm{WD}}$. The insets show the scalar potential at their respective position. The minimum of the potential, which corresponds to the realized field values, is marked in red. See text for details.
@@
We work in a weak coupling regime, where no scalar particles are admixed to the white dwarf matter; the whole effect of the scalar field comes from the presence of the expectation value. This can have an extremely strong backreaction on the white dwarf EOS. Firstly, let us focus on the scalar field coupled to electrons; we will comment on the axion model below. The model is described by two free parameters, the scalar mass $m_{\phi}$ and the coupling to electrons $d_{m_{e}}^{(2)}$, which we choose to be negative $d_{m_{e}}^{(2)}<0$. The former fixes the scalar field potential as
@@
up to relativistic corrections. At higher densities, $n>n_{c}$, $\phi$ will have a non-zero value and approach $\theta\to 1$ (where the electrons become massless) at large densities. Using the effective, density-dependent scalar field values, we can find the equation of state including the scalar field. Besides the scalar-field dependent electron mass, the scalar potential, which behaves like a vacuum energy, has to be included. To allow for a self-consistent addition of the scalar field effects, we use a simple model for the matter content. We describe the electrons as a free Fermi gas at finite temperature and combine this with one species of non-relativistic nuclei, following the ideal gas law. Consequently, after using charge neutrality, the EOS is given by
@@
When including the scalar field, this simple EOS shows a first-order phase-transition behavior (Fig. 2, top).
Figure 2: First-order phase transition behavior in the EOS. Top: Schematic sketch of the phase transition behavior, stable, metastable, and unstable branches, see text. Bottom: EOS in the presence of a quadratically coupled scalar field for different values of mass to coupling ratios $c_{m}$. The equation of state without a scalar field is shown in grey and the metastable branch in red. Note that for $c_{m}=0.01$, the phase transition happens at pressures and densities much lower than visible. The temperature is fixed at $T=10000~\mathrm{K}$ for this example.
@@
If, instead, the central density is smaller than the critical one ($\rho_{0}<\rho_{c}$), in principle, the stable (but sourced) EOS would be energetically favorable (if $\rho>\rho_{1}$). This would require a large increase in the scalar field, which is prevented by a potential barrier. Consequently, solutions with $\rho_{0}<\rho_{c}$ track the metastable branch of the EOS, resulting in an ordinary white dwarf.
@@
see also Fig. 2 (right), where we show the total pressure $p$ as a function of the mass density. These solutions are shown in various colors in the bottom panel of Fig. 2 for different values of $c_{m}$ and are energetically preferred to the SM white dwarf EOS, shown in thick grey.
For lower values of $c_{m}<0.0094$, the first-order phase transition behavior disappears. Again, there exists a metastable branch for densities $\rho<\rho_{c}$ with $\phi=0$ and a stable branch which now starts with $p(\rho)=0$ at non-zero densities $\rho=\rho_{2}$ and $\phi\neq 0$. The state for $\rho>\rho_{2}$ with sourced scalar is in a new ground state of matter (NGS) as shown in Ref. (Balkin et al., 2024).
Instead, for larger values $c_{m}>0.144$, the first-order phase transition is replaced by a second-order phase transition. This is because at large $c_{m}$ the pressure does not decrease at $\rho_{c}$, as is the case for the FOPT (see top panel of Fig. 2). Instead, it still increases, however, more slowly for $\rho>\rho_{c}$ than for $\rho<\rho_{c}$, and hence no Maxwell construction is needed. The effect ceases to exist with increasing $c_{m}$ as the slope of the pressure at $\rho>\rho_{c}$ tends towards the SM white dwarf equation of state. See also (Bartnick et al., 2025) for a discussion of the phase structure of this model as well as more details on the NGS.
@@
Instead of coupling to electrons, a new light scalar field might also couple to nucleons. This will typically lead to a new ground state of matter, resulting in drastic changes of the structure of white dwarfs. The process is therefore tightly constrained (see, for instance, Balkin et al. 2024, 2025). QCD axion models have such a coupling to nucleons but are a priori not testable in white dwarfs, since $n_{c}$ is too high. However, axion models, which still solve the strong CP problem but have a lower mass at a fixed decay constant, become testable as first discussed by (Balkin et al., 2024). Here we focus on the so-called $\mathbb{Z}_{\mathcal{N}}$-model for the lighter than usual QCD-axion developed by (Hook, 2018), where the axion mass is suppressed with a factor of roughly $(1/2)^{\mathcal{N}}$. For $\mathcal{N}=31$, in particular, this leads to a previously unconstrained cross-over phase-transition, which is different from the NGS, first- and second-order phase transition discussed above, but observable in white dwarfs. Due to its more complicated vacuum structure, there is also a metastable branch present. We discuss all details in App. A.2. Here we just show the resulting EOS (Fig. 3).
@@ 3 Calculation of models for white dwarfs
White dwarfs are the endproduct of the evolution for all stars with masses up to $\approx 8$ M, that is the vast majority of all stars. “Endproduct” means that the stars have no nuclear energy production and instead derive their energy output from gravitational contraction. Because of the extreme densities in the interior ($10^{6}$ g/cm3 and more) the electrons form a degenerate Fermi gas, almost independent of temperature, which dominates the EOS. In that case, gravitational contraction does not lead to heating, as in normal stars, but to a cooling process, which takes billions of years.
@@
The parameters used for a specific calculation are the total mass $M$ and total luminosity $L$. For the solution with a Runge-Kutta integration we need four boundary conditions. These are at the center $r(0)=0,l(0)=0$, and at the surface of the star $P(M)$ and $T(M)$. Because of this nature of the boundary conditions, the usual method to solve the equations is to assume numbers for the missing data at the center and at the surface, solve the equations from the surface to some fitting point, from the center to this fitting point, and iterate the guessed boundary values until both integrations agree at the fit point. In detail, $P(0),T(0)$ and $R(M)$ are guessed and iterated; from the assumed $R$, and the fixed parameters M and L the effective temperature and surface gravities can be calculated. These are then used to obtain $P(M)$ and $T(M)$ from a model of the stellar atmosphere with these parameters.
@@
In addition to the structure equations we also need an equation of state, and the opacity of the matter. In the outer envelope integration we use the H/He equation of state of (Saumon et al., 1995), augmented outside the range of their tables by our own version of the classical EOS. The latter includes the ideal gas for ions, partially degenerate electrons, and Coulomb interactions, which are calculated from numerical fits to Monte Carlo simulations of the One Component Plasma (DeWitt et al. (1996), as cited in (Chabrier & Potekhin, 1998)).
In the interior calculation (95% of the total mass) different versions of the EOS with scalar fields or with axions are used (see Sec. 2). The opacities are obtained from the OPAL tables (Iglesias & Rogers, 1996), and from the Los Alamos opacity tables (Colgan et al., 2016). Electron conduction data are from (Potekhin et al., 1999).
@@ 4 Mass-radius relation of white dwarfs
Numerous studies have discussed the theoretical relation between the mass and the radius of the white dwarfs, starting from the the zero-temperature models of (Chandrasekhar, 1935, 1939; Hamada & Salpeter, 1961) to the latest calculations including finite temperature effects of the Montreal  [^2] https://www.astro.umontreal.ca//~bergeron/CoolingModels/ and La Plata (Buenos Aires)  [^3] https://evolgroup.fcaglp.unlp.edu.ar/modelos.html groups (e.g. Bédard et al., 2017; Panei et al., 2000).
@@
Figure 4: Mass-radius relations with the classical SM equation of state for the effective temperatures $5000,15000,25000,35000,60000\,\mathrm{K}$. Color codes different temperature ranges; red: $5000-20000\,\mathrm{K}$, green: $20000-40000\,\mathrm{K}$, blue: $40000\,\mathrm{K}$. The crosses are the observed masses and radii described in the text. They use the same color coding as the theoretical relations, except for Sirius B, which we plot in magenta for better visibility. The dashed line is the $15000\,\mathrm{K}$ relation from the LaPlata group. The inset shows an enlarged plot around Sirius B.
@@
Figure 5: Mass-radius relations with Scalar equation of state for the effective temperatures $5000,15000,25000,35000$ and $60000\,\mathrm{K}$ (from bottom) and $c_{m}=0.03$. The five upper curves are stable branches for low masses, extended by metastable branches depending on the central density (see Fig. 2). The crosses are the observed masses and radii as in Fig. 4, which the theory with the new scalar field badly fails to explain. Coloring of objects and mass-radius relations as in Fig. 4.
@@
Figure 6: Dependence of the theoretical mass-radius relation for Teff = $25000\,\mathrm{K}$ on the parameter $c_{m}=0.01-0.14$. The objects in cyan have effective temperatures above 35000 K, Sirius B ($T_{\mathrm{eff}}=25369\pm 63\,\mathrm{K}$) is again shown in magenta. Note that the model with $c_{m}=0.01$ is very strongly first-order and thus shows the strongest deviation (see Fig. 2).
@@
Figure 7: Mass-radius relations with $\mathcal{N}=31,\mathbb{Z}_{\mathcal{N}}$-axion modified equation of state. The MR-relations from the stable (lower curves) and metastable (upper curves) branches of the EOS are shown for the effective temperatures $5000,15000,25000,35000$ and $60000\,\mathrm{K}$ (from the bottom). The metastable branches end, when the central density $n_{0}$ equals the critical one $n_{c}$. Increasing the central density beyond this point would trigger the phase transition in the star and lead to an unstable collapse, as denoted by the dark blue arrow. The metastable branch can explain nearly all well-observed white dwarfs, Sirius B, thought (shown in magenta), is in clear contradiction to the prediction of this model. Coloring of MR-relations and objects based on temperature as in Fig. 4.
@@
The dependence of the theoretical relation (MRR) on mass, luminosity, effective temperature, interior and outer chemical composition is well understood. The relation is routinely used to determine e.g. masses from spectroscopically determined surface gravities. The gravity $g=GM/R^{2}$ provides a second relation between mass and radius, such that both quantities can be determined.
@@
  • radius from surface gravity g + dynamical mass from white dwarfs in binary systems
  • radius and mass from total luminosity L + distance + effective temperature + surface gravity
@@
  • radius and mass from total luminosity L + distance + effective temperature + gravitational redshift
  • radius and mass from eclipsing white dwarfs in binary systems + gravitational redshift
Most of the methods involve an analysis of the spectra through comparison with theoretical models of the atmospheres to obtain effective temperatures and surface gravities. This typically involves errors of 3-5% for $T_{\text{eff}}$ and 15-20% for $g$. The distances - before the data from the astrometric satellite Gaia became available - could easily add another 20-30% error. Another problem is that the white dwarfs are strongly concentrated around an average mass of 0.6 solar masses and very few are near the low and high ends of the mass distribution. While the bulk of the MRR distribution did agree with the theoretical prediction, the shape was not empirically confirmed. All this contributed to the situation, that while the number of known white dwarfs had increased from 3 in 1920 to 359000 in 2021 (Gentile Fusillo et al., 2021), there were until recently only a handful of objects with significantly smaller errors than described above.
The situation changed dramatically with the data from eclipsing binaries with white dwarfs together with gravitational redshifts presented by (Parsons et al., 2017). This method is not only very accurate, but also independent of theoretical model atmospheres. For the first time a larger sample with typical errors of 2% for mass and radius became available.
@@
    a: (Parsons et al., 2017), b: (McGill et al., 2023), c: (Bond et al., 2017b), d: (Bond et al., 2017a), e: (Bond et al., 2015) and f: (Sahu et al., 2017)
As a first test of our approach, we calculate the MRR using the classical EOS described in Section 3 for masses between approximately 0.2 and 1.3 M for effective temperatures 5000, 15000, 25000, 35000, 60000 K, which covers the entire range of the observational sample.

### v2

# New ultralight scalar particles and the mass-radius relation of white dwarfs – the important role of Sirius B

@@ Abstract
astroparticle physics – white dwarfs – dark matter
pubyear: 2025pagerange: New ultralight scalar particles and the mass-radius relation of white dwarfs – the important role of Sirius BA.2
@@ 1 Introduction
Most prominently, there is the problem of dark matter, for which the SM has no candidate. Historically, dark matter searches focused on heavy, GeV-scale candidates. However, after decades of null results from indirect dark matter searches and direct detection experiments, particle theorists started to broaden the mass range they considered for their models. A particularly active field looks at ultralight dark matter (ULDM), also referred to as weakly interacting sub-eV particles (WISPs) (see e.g. Preskill et al., 1983; Abbott & Sikivie, 1983; Dine & Fischler, 1983; Arvanitaki et al., 2015). Instead of hiding dark matter at high energies or masses, detection is avoided by adding very light bosonic particles characterized by minuscule interactions with the Standard Model. In the end, this can be described by adding a new species of light scalar particles to the SM.
Another theoretical challenge for the Standard Model is the strong CP problem, which describes the violation of charge conjugation (C) and parity (P) symmetry: while the electro-weak interaction has an order one CP phase, no violation of CP symmetry has been detected in processes involving only the strong interaction. There is no explanation for this puzzle within the Standard Model, however, a very attractive solution is a new dynamical field, the bosonic axion (Peccei & Quinn, 1977b, a; Wilczek, 1978; Weinberg, 1978), which is also a viable dark matter candidate. Particles with similar properties, albeit not solving the strong CP problem, are also common in string theory and many other higher-dimensional models (Arvanitaki et al., 2010). These particles are commonly referred to as axion-like particles (ALPs).
@@
Astrophysics offers alternative approaches to constrain the properties of these scalar particles. The best-known example is based on stellar cooling: if additional light particles couple weakly to Standard Model matter, they get produced in stars and stellar remnants. Due to the weak coupling, they can easily escape and thus provide an additional energy loss channel for the star. Thus, stellar cooling in connection with stellar evolution data can be used to constrain these couplings (Raffelt, 1996; Springmann et al., 2025a). Considering, for example, QCD axions, generically, there are axion-nucleon couplings, which lead to axion production via nucleon bremsstrahlung. Requiring neutron stars not to cool faster than observed yields leading constraints on QCD axions (Buschmann et al., 2022; Springmann et al., 2025b).
@@
In this work, we will focus on a different approach to constraining particle physics from stars. Recently Hook & Huang (2018) have shown that new scalar fields with quadratic coupling to the Standard Model can develop a classical expectation value within stars - similarly to the Higgs mechanism, but localized within the star (see Fig. 1), a phenomenon that might occur for the QCD axion (Balkin et al., 2020). In scalar tensor theories of modified gravity, a similar phenomenon is known as scalarization (Damour & Esposito-Farese, 1993; Doneva et al., 2024; Ramazanoğlu & Pretorius, 2016; Staykov et al., 2018). This scalar field profile, especially the part that leaks out of the star can have drastic effects, see Hook & Huang (2018); Zhang et al. (2021) and Balkin et al. (2022, 2023). Moreover, the expectation value backreacts on the stellar structure and can lead to a new ground state (NGS) of stellar matter or a phase transition in the stellar equation of state (EOS). This would have drastic effects on astrophysical objects and observations, see for example Balkin et al. (2024, 2025); Gómez-Bañón et al. (2024); Kumamoto et al. (2025); Bartnick et al. (2025). Following this train of thought, in this work, we focus on two examples: a generic light scalar field coupled to electrons and leading to a first-order phase transition, and a lighter version of the QCD axion (Hook, 2018; Di Luzio et al., 2021; Banerjee et al., 2025b).
@@
We will use the observed white dwarf MRR to constrain these models of new physics. See also Crumpler et al. (2025) for an MRR constraint on linearly coupled ULDM models. We note that there has been a substantial recent improvement with the MRR observational data. While theoretically, within the SM, this relation is well understood and routinely used in the analysis of white dwarfs, the empirical evidence has been problematic until very recently (e.g. (Koester, 1987; Vauclair et al., 1997; Provencal et al., 1998; Joyce et al., 2018; Bédard et al., 2017)). The bulk of the white dwarf population agreed with the theoretical MRR, but the shape of the relation was almost unconstrained. This has changed dramatically over the last years due to the use of very large telescopes for high-quality optical spectra and orbital elements of eclipsing binaries, well calibrated ultraviolet (UV) spectra from the Hubble telescope, and in particular, excellent distance measurements from Gaia (Gaia Collaboration et al., 2023). This has now resulted in high-precision data for about 30 white dwarfs, which we use for the comparison with our theoretical models (see Table 1).
@@
In Section 2, we will outline the particle physics models and explain how they lead to a modified EOS. Section 3 details the calculation of white dwarf models, which lead to a (modified) MRR discussed in Section 4. The resulting limits on the particle physics models are shown and discussed in Section 5.
@@ 2 The equation of state with new light scalar fields
To derive constraints on new scalar fields, we must first understand how they influence the stellar structure of white dwarfs, which was first noted in Balkin et al. (2024) for light QCD axions. As detailed in the Appendix, we will focus on the so-called negligible-gradient limit where the length scale over which the scalar field changes is much smaller than the size of the star. Consequently, the effect of the scalar fields can be fully described by a modification of the white dwarf EOS. In this section, we will give a brief overview of the particle physics model and modified EOS. More details can be found in App. A and Balkin et al. (2024, 2025).
Figure 1: Schematic description of a scalar field profile emerging in a white dwarf. The main panel sketches the scalar field as a function of radius $r$ in and around a white dwarf. The dashed line marks the radius of the white dwarf $r=R_{\mathrm{WD}}$. The insets show the scalar potential at their respective positions. The minimum of the potential, which corresponds to the realized field values, is marked in red. See the text for details.
@@
We work in a weak coupling regime, where no scalar particles are admixed to the white dwarf matter; the whole effect of the scalar field comes from the presence of the expectation value. This can have an extremely strong backreaction on the white dwarf EOS. First, let us focus on the scalar field coupled to electrons; we will comment on the axion model below. The model is described by two free parameters, the scalar mass $m_{\phi}$ and the coupling to electrons $d_{m_{e}}^{(2)}$, which we choose to be negative $d_{m_{e}}^{(2)}<0$. The former fixes the scalar field potential as
@@
up to relativistic corrections. At higher densities, $n>n_{c}$, $\phi$ will have a non-zero value and approach $\theta\to 1$ (where the electrons become massless) at large densities. Using the effective, density-dependent scalar field values, we can find the EOS including the scalar field. Besides the scalar-field dependent electron mass, the scalar potential, which behaves like a vacuum energy, has to be included. To allow for a self-consistent addition of the scalar field effects, we use a simple model for the matter content. We describe the electrons as a free Fermi gas at finite temperature and combine this with one species of non-relativistic nuclei, following the ideal gas law. Consequently, after using charge neutrality, the EOS is given by
@@
When including the scalar field, this simple EOS shows a first-order phase-transition behaviour (Fig. 2, top).
Figure 2: First-order phase transition behaviour in the EOS. Top: Schematic sketch of the phase transition behaviour, stable, metastable, and unstable branches, see text. Bottom: EOS in the presence of a quadratically coupled scalar field for different values of mass to coupling ratios $c_{m}$. The EOS without a scalar field is shown in grey and the metastable branch in red. Note that for $c_{m}=0.01$, the phase transition happens at pressures and densities much lower than visible. The temperature is fixed at $T=10000~\mathrm{K}$ for this example.
@@
If, instead, the central density is smaller than the critical one ($\rho_{0}<\rho_{c}$), in principle, the stable (but sourced) EOS would be energetically favourable (if $\rho>\rho_{1}$). This would require a large increase in the scalar field, which is prevented by a potential barrier. Consequently, solutions with $\rho_{0}<\rho_{c}$ track the metastable branch of the EOS, resulting in an ordinary white dwarf.
@@
see also Fig. 2 (right), where we show the total pressure $p$ as a function of the mass density. These solutions are shown in various colours in the bottom panel of Fig. 2 for different values of $c_{m}$ and are energetically preferred to the SM white dwarf EOS, shown in thick grey.
For lower values of $c_{m}<0.0094$, the first-order phase transition behaviour disappears. Again, there exists a metastable branch for densities $\rho<\rho_{c}$ with $\phi=0$ and a stable branch which now starts with $p(\rho)=0$ at non-zero densities $\rho=\rho_{2}$ and $\phi\neq 0$. The state for $\rho>\rho_{2}$ with sourced scalar is in a new ground state of matter (NGS) as shown in Balkin et al. (2024).
Instead, for larger values $c_{m}>0.144$, the first-order phase transition is replaced by a second-order phase transition. This is because at large $c_{m}$ the pressure does not decrease at $\rho_{c}$, as is the case for the first-order phase transition (FOPT; see top panel of Fig. 2). Instead, it still increases, however, more slowly for $\rho>\rho_{c}$ than for $\rho<\rho_{c}$, and hence no Maxwell construction is needed. The effect ceases to exist with increasing $c_{m}$ as the slope of the pressure at $\rho>\rho_{c}$ tends towards the SM white dwarf equation of state. See also Bartnick et al. (2025) for a discussion of the phase structure of this model as well as more details on the NGS.
@@
Instead of coupling to electrons, a new light scalar field might also couple to nucleons. This will typically lead to a new ground state of matter, resulting in drastic changes of the structure of white dwarfs. The process is therefore tightly constrained (see, for instance, Balkin et al. 2024, 2025). QCD axion models have such a coupling to nucleons but are a priori not testable in white dwarfs, since $n_{c}$ is too high. However, axion models, which still solve the strong CP problem but have a lower mass at a fixed decay constant, become testable as first discussed by (Balkin et al., 2024). Here we focus on the so-called $\mathbb{Z}_{\mathcal{N}}$-model for the lighter than usual QCD-axion developed by Hook (2018), where the axion mass is suppressed with a factor of roughly $(1/2)^{\mathcal{N}}$. For $\mathcal{N}=31$, in particular, this leads to a previously unconstrained cross-over phase-transition, which is different from the NGS, first- and second-order phase transition discussed above, but observable in white dwarfs. Due to its more complicated vacuum structure, there is also a metastable branch present. We discuss all details in App. A.2. Here, we just show the resulting EOS (Fig. 3).
@@ 3 Calculation of models for white dwarfs
White dwarfs are the endproduct of the evolution for all stars with masses up to $\approx 8$ M, that is the vast majority of all stars. “Endproduct” means that the stars have no nuclear energy production and instead derive their energy output from gravitational contraction. Because of the extreme densities in the interior ($10^{6}$ g cm-3 and more) the electrons form a degenerate Fermi gas, almost independent of temperature, which dominates the EOS. In that case, gravitational contraction does not lead to heating, as in normal stars, but to a cooling process, which takes billions of years.
@@
The parameters used for a specific calculation are the total mass $M$ and total luminosity $L$. For the solution with a Runge-Kutta integration we need four boundary conditions. These are at the centre $r(0)=0,l(0)=0$, and at the surface of the star $P(M)$ and $T(M)$. Because of this nature of the boundary conditions, the usual method to solve the equations is to assume numbers for the missing data at the centre and at the surface, solve the equations from the surface to some fitting point, from the centre to this fitting point, and iterate the guessed boundary values until both integrations agree at the fit point. In detail, $P(0),T(0)$ and $R(M)$ are guessed and iterated; from the assumed $R$, and the fixed parameters M and L the effective temperature and surface gravities can be calculated. These are then used to obtain $P(M)$ and $T(M)$ from a model of the stellar atmosphere with these parameters.
@@
In addition to the structure equations we also need an equation of state, and the opacity of the matter. In the outer envelope integration we use the H/He equation of state of Saumon et al. (1995), augmented outside the range of their tables by our own version of the classical EOS. The latter includes the ideal gas for ions, partially degenerate electrons, and Coulomb interactions, which are calculated from numerical fits to Monte Carlo simulations of the One Component Plasma [DeWitt et al. (1996), as cited in Chabrier & Potekhin (1998)].
In the interior calculation (95% of the total mass) different versions of the EOS with scalar fields or with axions are used (see Sec. 2). The opacities are obtained from the OPAL tables (Iglesias & Rogers, 1996), and from the Los Alamos opacity tables (Colgan et al., 2016). Electron conduction data are from Potekhin et al. (1999).
@@ 4 Mass-radius relation of white dwarfs
Numerous studies have discussed the theoretical relation between the mass and the radius of the white dwarfs, starting from the zero-temperature models of Chandrasekhar (1935, 1939); Hamada & Salpeter (1961) to the latest calculations including finite temperature effects of the Montreal  [^2] https://www.astro.umontreal.ca//~bergeron/CoolingModels/ and La Plata (Buenos Aires)  [^3] https://evolgroup.fcaglp.unlp.edu.ar/modelos.html groups (e.g. Bédard et al., 2017; Panei et al., 2000).
@@
Figure 4: Mass-radius relations with the classical SM equation of state for the effective temperatures $5000,15\,000,25\,000,35\,000,60\,000\,\mathrm{K}$. Colour codes different temperature ranges; red: $5000-20000\,\mathrm{K}$, green: $20000-40000\,\mathrm{K}$, blue: $40000\,\mathrm{K}$. The crosses are the observed masses and radii described in the text. They use the same colour coding as the theoretical relations, except for Sirius B, which we plot in magenta for better visibility. The dashed line is the $15000\,\mathrm{K}$ relation from the LaPlata group. The inset shows an enlarged plot around Sirius B.
@@
Figure 5: Mass-radius relations with scalar EOS for the effective temperatures $5000,15000,25000,35000$ and $60000\,\mathrm{K}$ (from bottom) and $c_{m}=0.03$. The five upper curves are stable branches for low masses, extended by metastable branches depending on the central density (see Fig. 2). The crosses are the observed masses and radii as in Fig. 4, which the theory with the new scalar field badly fails to explain. Colouring of objects and mass-radius relations as in Fig. 4.
@@
Figure 6: Dependence of the theoretical mass-radius relation for $T_{\mathrm{eff}}$ = $25000\,\mathrm{K}$ on the parameter $c_{m}=0.01-0.14$. The objects in cyan have effective temperatures above 35000 K, Sirius B ($T_{\mathrm{eff}}=25369\pm 63\,\mathrm{K}$) is again shown in magenta. Note that the model with $c_{m}=0.01$ is very strongly first-order and thus shows the strongest deviation (see Fig. 2).
@@
Figure 7: Mass-radius relations with $\mathcal{N}=31,\mathbb{Z}_{\mathcal{N}}$-axion modified EOS. The MRRs from the stable (lower curves) and metastable (upper curves) branches of the EOS are shown for the effective temperatures 5000,15 000,25 000,35 0005000,15\,000,25\,000,35\,000, and $60\,000\,\mathrm{K}$ (from the bottom). The metastable branches end, when the central density $n_{0}$ equals the critical one $n_{c}$. Increasing the central density beyond this point would trigger the phase transition in the star and lead to an unstable collapse, as denoted by the dark blue arrow. The metastable branch can explain nearly all well-observed white dwarfs, Sirius B, thought (shown in magenta), is in clear contradiction to the prediction of this model. Colouring of MRRs and objects based on temperature as in Fig. 4.
@@
The dependence of the theoretical relation (MRR) on mass, luminosity, effective temperature, interior, and outer chemical composition is well understood. The relation is routinely used to determine for example masses from spectroscopically determined surface gravities. The gravity $g=GM/R^{2}$ provides a second relation between mass and radius, such that both quantities can be determined.
@@
  • radius from surface gravity g + dynamical mass from white dwarfs in binary systems,
  • radius and mass from total luminosity L + distance + effective temperature + surface gravity,
@@
  • radius and mass from total luminosity L + distance + effective temperature + gravitational redshift,
  • radius and mass from eclipsing white dwarfs in binary systems + gravitational redshift.
Most of the methods involve an analysis of the spectra through comparison with theoretical models of the atmospheres to obtain effective temperatures and surface gravities. This typically involves errors of 3-5% for $T_{\text{eff}}$ and 15-20% for $g$. The distances – before the data from the astrometric satellite Gaia became available – could easily add another 20-30% error. Another problem is that the white dwarfs are strongly concentrated around an average mass of 0.6 solar masses and very few are near the low and high ends of the mass distribution. While the bulk of the MRR distribution did agree with the theoretical prediction, the shape was not empirically confirmed. All this contributed to the situation, that while the number of known white dwarfs had increased from 3 in 1920 to 359000 in 2021 (Gentile Fusillo et al., 2021), there were until recently only a handful of objects with significantly smaller errors than described above.
The situation changed dramatically with the data from eclipsing binaries with white dwarfs together with gravitational redshifts presented by Parsons et al. (2017). This method is not only very accurate, but also independent of theoretical model atmospheres. For the first time a larger sample with typical errors of 2% for mass and radius became available.
@@
    a: Parsons et al. (2017), b: McGill et al. (2023), c: Bond et al. (2017b), d: Bond et al. (2017a), e: Bond et al. (2015) and f: Sahu et al. (2017)
As a first test of our approach, we calculate the MRR using the classical EOS described in Section 3 for masses between approximately 0.2 and 1.3 M for effective temperatures 5000, 15 000, 25 000, 35 000, and 60 000 K, which covers the entire range of the observational sample.